OpenStudy (tux):

A remote controlled car is moving in a vacant parking lot. The velocity of the car as a function of time is given by  v= [5.00 m/s – (0.0180 m/s2)t^2 ]i+[2.00 m/s + (0.550 m/s2)t ]j . a) What are ax(t) and ay(t), the x- and y- components of the velocity of the car as functions ? b) What are the magnitude and direction of the velocity of the car at t= 8.00 s?

OpenStudy (mos1635):

Vx=5-0.018t^2 Vy=2+0.55t ax=dVx/dt ax=-0.036t ay=dVy/dt ay=0.55